A car of mass 100 kg starts moving on a horizontal road from a point o take the horizontal plane containing o as reference level for gravitational potential energy g=10m/s^2 the car reaches point A with speed 72km/hr calculate the mechanical energy
1
Expert's answer
2020-10-19T13:24:04-0400
"E=0.5mv^2\\E=0.5(100)\frac{72^2}{3.6^2}=20000\ J"